# Community detection in the hypergraph stochastic block model and reconstruction on hypertrees

### v1

@@ A.4 Fine estimates
Appendix B HSBM weak recovery upper bound
We say an event happens a.a.s. (asymptotically almost surely) if it happens with probability $1-o(1)$ as $n\to\infty$.
@@ Appendix B HSBM weak recovery upper bound
We first describe our algorithm. Suppose we are given a HSBM instance $G=(V,E)$. We say a partition $Y:V\to\{\pm\}$ is almost balanced if |#$\left|\#Y^{-1}(+)-\frac{n}{2}\right|\leq n^{0.6}$. We say a hyperedge $e\in E$ is in-community with respect to $Y$ if $Y_{v}$s are equal for all $v\in e$. We say a partition $Y:V\to\{\pm\}$ is good if the number of in-community hyperedges is $\left[\frac{an}{r}-n^{0.6},\frac{an}{r}+n^{0.6}\right]$, and bad otherwise. The algorithm looks for a good almost balanced partition. If such a partition exists, return any of them; otherwise report failure.
We say a partition $Y$ is correlated with $X$ (the true partition) if $d_{H}(Y,X)\leq\left(\frac{1}{2}-\epsilon\right)n$, and uncorrelated otherwise. We prove that for some $\epsilon>0$, a.a.s. the following are true.
  1. (a)
    There exists a good almost balanced partition.
  2. (b)
    All almost balanced partitions that are uncorrelated with $X$ are bad.
If (a) holds, our algorithm returns a good partition. If, in addition, (b) holds, the returned partition is correlated with the true partition $X$, and the algorithm succeeds. It remains to prove (a) and (b).
Proof of (a)
We show that $X$ is a.a.s. a good almost balanced partition. By central limit theorem, $X$ is a.a.s. almost balanced. For such $X$, the number of in-community sets $S\in\binom{V}{r}$ is $\binom{|X|}{r}+\binom{n-|X|}{r}=(1\pm O(n^{-1}))\frac{n^{r}}{2^{r-1}\cdot r!}$. Each such set is an hyperedge independently with probability $\frac{a}{\binom{n}{r-1}}=(1\pm O(n^{-1}))\frac{a(r-1)!}{n^{r-1}}$. By Chebyshev inequality, a.a.s. the number of in-community hyperedges is $\left[\frac{an}{r}-n^{0.6},\frac{an}{r}+n^{0.6}\right]$.
Proof of (b)
  1. 1.
    $X$ is almost balanced.
  2. 2.
    Total number of hyperedges $m$ is in $\left[\frac{dn}{r}-n^{0.6},\frac{dn}{r}-n^{0.6}\right]$.
We show that (b) is true conditioned on $m$. We consider the following random hypergraph model $G(n,m,X)$. For each of the $m$ edges, generate a sequence $(x_{1},\ldots,x_{r})$ of $r$ bits according to the following distribution
$p_{x_{1},\ldots,x_{r}}=\left\{\begin{array}[]{ll}\frac{1}{2}\cdot\frac{a}{a-b+2^{r-1}b}&\text{if}~{}x_{1}=\cdots=x_{r},\\ \frac{1}{2}\cdot\frac{b}{a-b+2^{r-1}b}&\text{o.w.}\end{array}\right.$ (219)
Then choose $r$ vertices $v_{1},\ldots,v_{r}$ independently according to $v_{i}\sim\operatorname{\mathrm{Unif}}(X^{-1}(x_{i}))$. Finally, add edge $e=(v_{1},\ldots,v_{r})$ to the hypergraph.
$c_{++}=\frac{1}{4}+\frac{\delta}{2}\pm O(n^{-0.4}),$ $c_{+-}=\frac{1}{4}-\frac{\delta}{2}\pm O(n^{-0.4}),$ (220)
$c_{-+}=\frac{1}{4}-\frac{\delta}{2}\pm O(n^{-0.4}),$ $c_{--}=\frac{1}{4}+\frac{\delta}{2}\pm O(n^{-0.4}).$
[…]
$\mathbb{P}[x^{\prime}_{i}=+|x_{i}=+]=\frac{1}{2}+\delta\pm O(n^{-0.4}),\qquad\mathbb{P}[x^{\prime}_{i}=+|x_{i}=-]=\frac{1}{2}-\delta\pm O(n^{-0.4}).$ (221)
Therefore, under $G(n,m,X)$, each edge is in-community with respect to $Y$ with probability
$~{}\sum_{x_{1},\ldots,x_{r}\in\{\pm\}}p_{x_{1},\ldots,x_{r}}\left(\prod_{i\in[r]}\mathbb{P}[x^{\prime}_{i}=+|x_{i}]+\prod_{i\in[r]}\mathbb{P}[x^{\prime}_{i}=-|x_{i}]\right)$ (222)
[…]
By Eq. (5), for $\epsilon>0$ close enough to 00, the last expression is $\exp\left(-n(\log 2+\epsilon^{\prime})\pm o(n)\right)$ for some $\epsilon^{\prime}>0$. Therefore every uncorrelated balanced partition $Y$ is good with probability $\exp\left(-n(\log 2+\epsilon^{\prime})\pm o(n)\right)$. On the other hand, the total number of partitions is $2^{n}$. By union bound, a.a.s. there are no uncorrelated balanced partitions. This finishes the proof of (b). ∎
@@
In this proof only, define $f_{r}(\lambda)=d_{\operatorname{\mathrm{KL}}}\left(\lambda+\frac{1-\lambda}{2^{r-1}}\left\|\frac{1}{2^{r-1}}\right.\right)$. Computation shows that
Therefore, for any $\delta>0$, there exists $\lambda_{0}=\lambda_{0}(r,\delta)>0$ such that for all $|\lambda|\leq\lambda_{0}$, we have $f_{r}(\lambda)>\left(\frac{2^{r-1}-1}{2}-\delta\right)\lambda^{2}$. Furthermore, because $f_{r}(\lambda)$ is monotone increasing for $\lambda>0$ and monotone decreasing for $\lambda<0$, we can take $\lambda_{0}$ such that $f_{r}(\lambda)>\left(\frac{2^{r-1}-1}{2}-\delta\right)\min\{\lambda^{2},\lambda_{0}^{2}\}$ for all $\lambda\in\left[-\frac{1}{2^{r-1}-1},1\right]$.
@@ Appendix C Discussions
C.1 Comparison with $q$-community SBM
The HSBM is a generalization of the two-community SBM to hypergraphs. Another natural generalization is the $q$-community SBM. In the $q$-SBM, $n$ vertices a given i.i.d. $\operatorname{\mathrm{Unif}}([q])$ labels $X_{v}$. An edge $(u,v)$ is added with probability $\frac{a}{n}$ if $X_{u}=X_{v}$, and with probability $\frac{b}{n}$ if $X_{u}\neq X_{v}$. The goal of the weak recovery problem is to recover a non-trivial fraction of the labels given the generated graph.
For the $q$-SBM, we define
@@ C.1 Comparison with $q$-community SBM
$q=4$, $\lambda<0$ (not tight) not tight tight
$q\geq 5$, $\lambda>0$ (not tight) not tight not tight
$q\geq 5$, $\lambda<0$ (not tight) not tight not tight
Table 2: Tightness of KS threshold for symmetric $q$-SBM weak recovery. An entry in parentheses means it is implied by another entry in the same row. References: Possibility above KS: [AS15]. Tightness for $q\leq 4$, large $d$: [MSS23]. Non-tightness for $q=4$, $\lambda<0$, small $d$ and $q\geq 5$: [AS15] (see also [BMNN16]).
Our current knowledge about tightness of the KS threshold for $q$-SBM weak recovery is summarized in Table 2. Comparing Table 1 and Table 2, we make a few interesting observations. In both models, in the large $d$ regime, as $r$ and $q$ goes larger, the KS threshold goes from tight to non-tight. However, the HSBM has a small $d$ regime ($\lambda\geq\frac{1}{5}$) where KS is tight for any $r$, but the $q$-SBM lacks such a regime.
For the $4$-SBM, it is conjectured that tightness of KS is different for $\lambda>0$ and $\lambda<0$ cases. It is known that for $\lambda<0$ and small $d$, KS is not tight, and conjectured that for $\lambda>0$, KS is always tight. We do not have reason to expect similar behavior exists in the HSBM.
We remark that Table 2 is also our current knowledge about tightness of the KS threshold for the $q$-ary Potts model on a tree. Tightness for $q=3$, large $d$ and non-tightness for $q\geq 5$ was established in [Sly11], and the results for $q=4$ was by [MSS23].
@@ C.2 Methods for reconstruction on trees and hypertrees
For the upper bound (reconstruction), the Kesten-Stigum bound is simple and works universally. When it is not tight, upper bounds are proved by establishing non-contraction properties. Suppose for a certain non-trivial channel $P$, we have $\operatorname{\mathrm{BP}}(P)\geq P$, where $\geq$ is a certain channel preorder that respects the $\operatorname{\mathrm{BP}}$ operator (which is often the degradation preorder), then by iterating we get $\operatorname{\mathrm{BP}}^{k}(\operatorname{\mathrm{Id}})\geq P$ for all $k\geq 0$, and reconstruction is possible. [Sly09] used freezing probability as the information measure to prove reconstruction results for the random coloring model, which can be viewed as taking $\geq$ to be the degradation preorder and $P$ to be an erasure channel. [Sly11, MSS23] used $\chi^{2}$-capacity (also called magnetization) to prove reconstruction results for the Potts model in the large degree regime (via the large degree asymptotics method). This was also used in [LN19] for the asymmetric Ising model and in [GP23b] for the BOHT model. This work (Theorem 7) also uses this method. Sometimes the channel $P$ is designed explicitly, such as [SZ16, Zha17] which established lower order terms of the reconstruction threshold for the random coloring model and NAE-$k$-SAT model (which is BOHT model with $\lambda=-\frac{1}{2^{r-1}-1}$).
For the lower bound (non-reconstruction), there are two particularly successful methods. One method proves contraction of certain information measures by using strong data processing inequalities (SDPIs) and subadditivity. [BCMR06] proved tightness of KS threshold for roughly symmetric Ising models by using $\chi^{2}$-information. [KF09] proved non-reconstruction results for Potts model using this method with symmetric KL (SKL) information, and [GP23a] obtained improved results by using mutual information. [GP23b] proved non-reconstruction results for the BOHT model by using a multi-terminal generalization of the method with SKL information and $\chi^{2}$-information. The advantage of this method is that it is often simple to obtain, works for any degree, and can be easily generalized to different tree structures (non-Galton-Watson trees). However, the choice of the information measure is mysterious and it is unclear what is the best information measure to use.
The other method is Sly’s method, which uses large degree asymptotics and robust reconstruction. This method was introduced in [Sly09, Sly11] for the random coloring model and Potts model. [LN19] used the method for the asymmetric Ising model. [MSS23] improved this method by analyzing more orders in the Taylor expansion, and removing the need for a certain concentration result. This work (Theorem 5) uses the hypertree verion of Sly’s method. The advantage of this method is that it often gives very tight results. The disadvantage is that it needs a large degree assumption and requires heavy computation.
@@ C.3 Further directions
Theorem 2 gives asymptotic bounds on the weak recovery threshold, where there is a factor $\Theta(r)$ gap. In [BMNN16], similar asymptotic bounds were obtained for the $q$-SBM, and there the gap is constant factor. Reducing the gaps in the asymptotic bounds is an interesting question.
For $r=5,6$ and $\lambda>0$, we know that the KS threshold is tight for small $d$ and large $d$. If there is a certain kind of convexity property, then we could conclude that KS is tight for all $d$.
For $r\geq 7$ and $\lambda>0$, we know that the KS threshold is tight for small $d$ and not tight for large $d$. Therefore, it could be possible such that there is a $d^{*}$ for which KS is tight for $d<d^{*}$ and not tight for $d>d^{*}$. It is not even clear that such a $d^{*}$ exists.

### v2

@@ A.4 Fine estimates
Appendix B Proofs of remaining results
In this section we prove our remaining main results, Corollary 3 (non-tightness of the KS threshold), Theorem 8 (asymptotic BOHT reconstruction threshold) and Theorem 4 (asymptotic HSBM weak recovery bounds).
B.1 Non-tightness of the KS threshold
Proof of Corollary 3.
For $\lambda\in\left[-\frac{1}{2^{r-1}-1},1\right]$, define
$f_{r}(\lambda)=d_{\operatorname{\mathrm{KL}}}\left(\lambda+\frac{1-\lambda}{2^{r-1}}\left\|\frac{1}{2^{r-1}}\right.\right).$ (220)
By Theorem 2, if for some $(r,\lambda)$ we have
$f_{r}(\lambda)>r(r-1)\lambda^{2}\log 2,$ (221)
then for this $(r,\lambda)$ the KS threshold is not tight for HSBM weak recovery.
For $r=5,6$, we take $\lambda_{0}$ as in Table 1 and verify that Eq. (221) holds for all $\lambda\in\left[-\frac{1}{2^{r-1}-1},\lambda_{0}\right)$. Therefore the KS threshold is not tight for such $(r,\lambda)$.
It remains to prove that for all $r\geq 7$, there exists $\lambda_{0}(r)>0$ such that Eq. (221) holds for all $\lambda\in\left[-\frac{1}{2^{r-1}-1},\lambda_{0}\right)$. For $\lambda\in\left[-\frac{1}{2^{r-1}-1},0\right)$, we apply Lemma 19 with $x=\lambda+\frac{1-\lambda}{2^{r-1}}$, $y=\frac{1}{2^{r-1}}$ and get
$f_{r}(\lambda)\geq\frac{2^{r-1}-1}{2}\lambda^{2}>r(r-1)\lambda^{2}\log 2$ (222)
where the last step holds for all $r\geq 7$.
For $\lambda>0$, we observe that
So for $r\geq 7$, there exists $\lambda_{0}=\lambda_{0}(r)>0$ such that Eq. (221) holds for all $\lambda\in\left(0,\lambda_{0}\right)$. Combining both cases we finish the proof. ∎
Lemma 19.
For $0\leq x\leq y\leq\frac{1}{2}$, we have
$d_{\operatorname{\mathrm{KL}}}\left(x\|y\right)\geq\frac{(x-y)^{2}}{2y(1-y)}.$ (224)
Proof.
We first consider the case $x>0$. Let $t=y-x$. Then
$d_{\operatorname{\mathrm{KL}}}\left(x\|y\right)$ $=x\log\left(1-\frac{t}{y}\right)+(1-x)\log\left(1+\frac{t}{1-y}\right)$ (225)
$=(-y+t)\sum_{k\geq 1}\frac{1}{k}\left(\frac{t}{y}\right)^{k}+(1-y+t)\sum_{k\geq 1}(-1)^{k-1}\frac{1}{k}\left(\frac{t}{1-y}\right)^{k}$
$=\left(-t+\sum_{k\geq 2}\frac{t^{k}y^{-k+1}}{k(k-1)}\right)+\left(t+\sum_{k\geq 2}(-1)^{k}\frac{t^{k}(1-y)^{-k+1}}{k(k-1)}\right)$
$=\sum_{k\geq 2}\frac{t^{k}}{k(k-1)}\left(y^{-k+1}+(-1)^{k}(1-y)^{-k+1}\right).$
Because $y\leq\frac{1}{2}$, the summand is non-negative for all $k\geq 2$. Taking $k=2$, we get
$d_{\operatorname{\mathrm{KL}}}\left(x\|y\right)\geq\frac{1}{2}t^{2}\left(y^{-1}+(1-y)^{-1}\right)=\frac{(y-x)^{2}}{2y(1-y)}.$ (226)
The case $x=0$ follows by continuity. ∎
B.2 Asymptotic BOHT reconstruction threshold
@@
Upper bound. Let $w=w^{*}+\epsilon$. By Lemma 15, it suffices to prove for $(r-1)d\lambda^{2}=w$. By Lemma 10(i) and definition of $w^{*}$, there exists $x_{0}\in(0,1]$ such that $g_{r,w}(x_{0})>x_{0}$. Let $\epsilon=g_{r,w}(x_{0})-x_{0}$. Take $d_{0}=d_{0}(r,\epsilon)$ in Prop. 9. By Prop. 9, for $d\geq d_{0}$ and $(r-1)d\lambda^{2}=w$, for any BMS channel $P$ with $C_{\chi^{2}}(P)\geq x_{0}$, we have
Lower bound. Let $w=w^{*}-\epsilon$. By Lemma 10(i) and definition of $w^{*}$, we have $g_{r,w}(x)<x$ for all $x\in(0,1]$. By the same proof as Prop. 6, for any $\epsilon^{\prime}>0$ there exists $d_{0}=d_{0}(r,\epsilon^{\prime})$ such that $\lim_{k\to\infty}C_{\chi^{2}}(M_{k})<\epsilon^{\prime}$. The rest of the proof is the same as the proof of Theorem 5. ∎
B.3 Asymptotic HSBM weak recovery bounds
Proof of Theorem 4(ii).
is a polynomial with non-negative coefficients. Therefore for $0\leq x\leq 1$, $\frac{s_{r,w}(x)}{x}$ is maximized at $x=1$, taking value $w2^{r-2}=1$. So $w^{*}(r)\geq w$. Applying Theorem 8 and [Gu23, Theorem 5.15] we finish the proof. ∎
@@ B.3 Asymptotic HSBM weak recovery bounds
Recall function $f_{r}$ defined in Eq. (220). By Eq. (223), for any $\delta>0$, there exists $\lambda_{0}=\lambda_{0}(r,\delta)>0$ such that for all $|\lambda|\leq\lambda_{0}$, we have $f_{r}(\lambda)>\left(\frac{2^{r-1}-1}{2}-\delta\right)\lambda^{2}$. Furthermore, because $f_{r}(\lambda)$ is monotone increasing for $\lambda>0$ and monotone decreasing for $\lambda<0$, we can take $\lambda_{0}$ such that $f_{r}(\lambda)>\left(\frac{2^{r-1}-1}{2}-\delta\right)\min\{\lambda^{2},\lambda_{0}^{2}\}$ for all $\lambda\in\left[-\frac{1}{2^{r-1}-1},1\right]$.
@@ Appendix C Discussions
C.1 Comparison with the $q$-community SBM
The HSBM is a generalization of the two-community SBM to hypergraphs. Another natural generalization is the $q$-community SBM ($q$-SBM). In this model, $n$ vertices are given i.i.d. $\operatorname{\mathrm{Unif}}([q])$ labels $X_{v}$. An edge $(u,v)$ is added with probability $\frac{a}{n}$ if $X_{u}=X_{v}$, and with probability $\frac{b}{n}$ if $X_{u}\neq X_{v}$. The goal of the weak recovery problem is to recover a non-trivial fraction of the labels given the generated graph.
For the $q$-SBM, define
@@ C.1 Comparison with the $q$-community SBM
$q=4$, $\lambda<0$ unknown not tight tight
$5\leq q\leq 10$, $\lambda>0$ unknown unknown unknown
$5\leq q\leq 10$, $\lambda<0$ unknown not tight unknown
$q\geq 11$, $\lambda>0$ unknown unknown not tight
$q\geq 11$, $\lambda<0$ never tight (not tight) (not tight)
Table 3: Tightness of the KS threshold for symmetric $q$-SBM weak recovery. An entry in parentheses means it is implied by another entry in the same row. References: Possibility above KS: [AS15]. Tightness results: [MSS23]. Non-tightness results: [AS15] (all non-tightness results) and [BMNN16] (for $q\geq 5$).
Our current knowledge about tightness of the KS threshold for $q$-SBM weak recovery is summarized in Table 3. Comparing Table 2 and Table 3, we make a few interesting observations. In both models, in the large $d$ regime, as $r$ and $q$ goes larger, the KS threshold goes from tight to non-tight. However, the HSBM has a small $d$ regime ($\lambda\geq\frac{1}{5}$) where KS is tight for any $r$, but the $q$-SBM is not known to have such a regime.
For the $4$-SBM, it is conjectured that tightness of the KS threshold is different for $\lambda>0$ and $\lambda<0$ cases. It is known that for $\lambda<0$ and small $d$, KS is not tight, and conjectured that for $\lambda>0$, KS is always tight. We expect similar behavior to hold for the HSBM with $r=5,6$.
Unlike the HSBM case, we know more about tightness of the KS threshold for $q$-ary Potts model on a tree than the $q$-SBM. For $q\geq 5$, it is known that the KS threshold is never tight ([Sly11, MSS23]) for the Potts model. However, as shown in Table 3, there are several regimes with $q\geq 5$ where the KS threshold is not known to be not tight for the $q$-SBM.
@@ C.2 Methods for reconstruction on trees and hypertrees
For the upper bound (reconstruction), the Kesten-Stigum bound is simple and works universally. When it is not tight, upper bounds are proved by establishing non-contraction properties. Suppose for a certain non-trivial channel $P$, we have $\operatorname{\mathrm{BP}}(P)\geq P$, where $\geq$ is a certain channel preorder that respects the $\operatorname{\mathrm{BP}}$ operator (which is often the degradation preorder), then by iterating we get $\operatorname{\mathrm{BP}}^{k}(\operatorname{\mathrm{Id}})\geq P$ for all $k\geq 0$, and reconstruction is possible. [Sly09] used freezing probability as the information measure to prove reconstruction results for the random coloring model, which can be viewed as taking $\geq$ to be the degradation preorder and $P$ to be an erasure channel. [Sly11, MSS23] used $\chi^{2}$-capacity (also called magnetization) to prove reconstruction results for the Potts model in the large degree regime (via the large degree asymptotics method). This was also used in [LN19] for the asymmetric Ising model and in [GP23b] for the BOHT model. The current work uses this method in Theorem 8. Sometimes the channel $P$ is designed explicitly, such as [SZ16, Zha17] which established lower order terms of the reconstruction threshold for the random coloring model and the NAE-$k$-SAT model (in other words, the BOHT model with $\lambda=-\frac{1}{2^{r-1}-1}$).
For the lower bound (non-reconstruction), there are two particularly successful methods. One method proves contraction of certain information measures using strong data processing inequalities (SDPIs) and subadditivity. [BCMR06] proved tightness of the KS threshold for roughly symmetric Ising models using $\chi^{2}$-information. [KF09] proved non-reconstruction results for Potts model using this method with symmetric KL (SKL) information, and [GP23a] obtained improved results using mutual information. [GP23b] proved non-reconstruction results for the BOHT model using a multi-terminal generalization of the method with SKL information and $\chi^{2}$-information. The advantage of this method is that it is often simple to obtain, works for any degree, and can be easily generalized to different tree structures (non-Galton-Watson trees). However, the choice of the information measure is mysterious and it is unclear what is the best information measure to use.
The other method is Sly’s method, which uses large degree asymptotics and robust reconstruction. This method was introduced in [Sly09, Sly11] for the random coloring model and the Potts model. [LN19] used the method for the asymmetric Ising model. [MSS23] improved this method by analyzing more orders in the Taylor expansion, and removing the need for a certain concentration result. The current work uses a hypertree verion of Sly’s method in Theorem 5. The advantage of this method is that it often gives tight results. The disadvantage is that it needs a large degree assumption and requires heavy computation.
@@ C.3 Further directions
For $r=5,6$ and $\lambda>0$, Table 2 shows that the KS threshold is tight for small $d$ and large $d$. If there is a certain kind of convexity property, then we could conclude that KS is tight for all $d$.
For $r\geq 7$ and $\lambda>0$, Table 2 shows that the KS threshold is tight for small $d$ and not tight for large $d$. Therefore, it seems possible that there exists $d^{*}=d^{*}(r)$ for which KS is tight for $d<d^{*}$ and not tight for $d>d^{*}$.
Asymptotic HSBM weak recovery bounds
Theorem 4 gives asymptotic bounds on the weak recovery threshold, where there is a factor $\Theta(r)$ gap. In [BMNN16], similar asymptotic bounds were obtained for the $q$-SBM, and the gap there is of a factor of two. Reducing the gaps in the asymptotic bounds for HSBM is an interesting question. Evidences suggest that the upper bound (Theorem 4(i)) might be closer to the true threshold as $r\to\infty$. The lower bound (Theorem 4(i)) could be not tight for two reasons. One is that the HSBM weak recovery threshold may be different from the BOHT reconstruction threhsold; the other is that the constant $\frac{1}{2^{r-2}}$ is not an asymptotically tight bound of $w^{*}(r)$ in Eq. (23). We choose $\frac{1}{2^{r-2}}$ so that the statement of Theorem 4(i) works for any $r\geq 3$ rather than as $r\to\infty$.
Non-uniform HSBM
An interesting direction of generalization is the non-uniform HSBMs, where hyperedges may have different sizes. This model has a sequence of non-negative real parameters $(a_{r},b_{r})_{r\geq 2}$ with finitely many non-zero entries (not depending on $n$, the number of vertices). First, the vertices are given i.i.d. $\operatorname{\mathrm{Unif}}(\{\pm\})$ labels. Then each set $S$ of vertices with $|S|\geq 2$ is added as an hyperedge with probability $\frac{a_{|S|}}{\binom{n}{|S|-1}}$ or $\frac{b_{|S|}}{\binom{n}{|S|-1}}$ depending on whether the set $S$ is monochromatic. The KS threshold is defined as $\sum_{r\geq 2}(r-1)d_{r}\lambda_{r}^{2}=1$ where $d_{r}$ and $\lambda_{r}$ are defined from $a_{r}$ and $b_{r}$ as in Eq. (1). [CEH23] conjectured that weak recovery is possible above the KS threshold, and partial progress was made in [DWZ21]. By analyzing the corresponding non-uniform BOHT model, one could show that weak recovery is impossible below the KS threshold under certain conditions. For example, by adapting the multi-terminal SDPI method from [GP23b], one can show that if $a_{r}\geq b_{r}$ for $r\leq 4$ and $a_{r}=b_{r}=0$ for $r\geq 5$, then weak recovery is impossible below the KS threshold. Furthermore, our robust reconstruction result (Theorem 7) seems to still hold in the non-uniform case. On the other hand, by adapting our proof of Theorem 2 one could prove that weak recovery is possible whenever
$\sum_{r\geq 2}\frac{d_{r}}{r}\cdot d_{\operatorname{\mathrm{KL}}}\left(\lambda_{r}+\frac{1-\lambda_{r}}{2^{r-1}}\left\|\frac{1}{2^{r-1}}\right.\right)>\log 2.$ (231)